A piece of copper metal of mass 6.5kg is heated from 20.5°C to 324.3°C calculate the heat absorbed in kilo calories by the metal specific heat capacity of CU is 0.385 j/g°C


1
Expert's answer
2021-11-05T00:56:35-0400

"\Delta\>H=mc\Delta\>T"


"6.5kg=6.5\u00d71000=6500g"

"\Delta\>T=324.3-20.5=303.8°C"

"\Delta\>H=6500×0.385×303.8"

"=760259.5 \>Joules"

"1kilocalorie=4184Joules"


"Heat\> absorbed= \frac{760259.5}{4184}"

"=181.7 \>kilocalories"